OpenStudy (anonymous):

what is general solution for tan(θ/4)=1?

jimthompson5910 (jim_thompson5910):

do you agree that tan(x) = sin(x)/cos(x) ??

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so that means if tan(theta) = 1, then sin(theta) = cos(theta)

jimthompson5910 (jim_thompson5910):

look on the unit circle for points that have the same x and y coordinates

OpenStudy (anonymous):

do you mean like (√2/2, √2,2)

jimthompson5910 (jim_thompson5910):

yes, what is the angle that corresponds to that point?

OpenStudy (anonymous):

45

jimthompson5910 (jim_thompson5910):

good, or pi/4 radians

jimthompson5910 (jim_thompson5910):

where else?

jimthompson5910 (jim_thompson5910):

hint: look in Q3

OpenStudy (anonymous):

5pi/4

OpenStudy (anonymous):

so would those be for the two equations

jimthompson5910 (jim_thompson5910):

so if theta = pi/4 or theta = 5pi/4, then tan(theta) = 1

jimthompson5910 (jim_thompson5910):

now these aren't the only 2 solutions. There are infinitely more

jimthompson5910 (jim_thompson5910):

we have to account for all of the coterminal angles

jimthompson5910 (jim_thompson5910):

eg: theta = pi/4 add on 2pi (pi/4) + 2pi = pi/4 + 8pi/4 = 9pi/4 in general, we just write \[\Large \frac{\pi}{4} + 2\pi*k\] where k is any integer (whole number)

jimthompson5910 (jim_thompson5910):

we do the same for 5pi/4 to get \[\Large \frac{5\pi}{4} + 2\pi*k\]

jimthompson5910 (jim_thompson5910):

with me so far?

OpenStudy (anonymous):

yes this is making much more sense

jimthompson5910 (jim_thompson5910):

ok so the theta inside the tangent is actually theta/4 which means ------------------- if \[\Large \tan\left(\frac{\theta}{4}\right) = 1\] then \[\Large \frac{\theta}{4} = \frac{\pi}{4}+2\pi*k\] or \[\Large \frac{\theta}{4} = \frac{5\pi}{4}+2\pi*k\] agreed?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

the next step is to multiply both sides by 4 to completely isolate theta So let's do that to the first equation \[\Large \frac{\theta}{4} = \frac{\pi}{4}+2\pi*k\] \[\Large 4*\frac{\theta}{4} = 4*\left(\frac{\pi}{4}+2\pi*k\right)\] \[\Large \theta = \pi+8\pi*k\] and then to the other equation as well \[\Large \frac{\theta}{4} = \frac{5\pi}{4}+2\pi*k\] \[\Large 4*\frac{\theta}{4} = 4*\left(\frac{5\pi}{4}+2\pi*k\right)\] \[\Large \theta = 5\pi+8\pi*k\] ---------------------- So, in the end, we get \[\Large \theta = \pi+8\pi*k\] or \[\Large \theta = 5\pi+8\pi*k\]

OpenStudy (anonymous):

wow thanks so much!!

jimthompson5910 (jim_thompson5910):

you're welcome